OpenStudy (anonymous):

Hi, Can anyone help me with this question plzz A 2.0*10^2 g pendulum bob is raised 22cm above its rest position. The bob is is released, and it reaches its maximum speed as it passes the rest position. a) Calculate its maximum speed at that point by applying the law of conservation of energy and assuming that the efficiency is 100%. b) Repeat (a) if the efficiency is 94%.

OpenStudy (anonymous):

 Eqn's to use: E=U+K=mgh+0.5mv^2 U=potential energy K=Kinetic energy E=Mechanical energy At any given time the system has a constant E therefore the E at the top of the swing equals the E at the bottom. At the top the pendulum has no velocity so K=0 and E=mgh At the bottom U=0 so EM=0.5mv^2 For part 2 the E at the bottom is 0.94E at the top